OpenStudy (anonymous):

The number N of bacteria after t hours is N(t)=A \cdot 2^t. Initially there are 50 bacteria. How many are there after 10 hours?.

OpenStudy (anonymous):

Can you rewrite your equation using the equation tool please? I don't understand what your equation is.

OpenStudy (michele_laino):

from the text of your problem, we can write: \[N(t)=50*2^{t}\] where t is time, so you, please have to substitute t =10 into the above equation

OpenStudy (michele_laino):

@tom982 what do you think about that, please?

OpenStudy (anonymous):

Looks good @Michele_Laino, I had no idea \cdot was a LaTeX character, I thought it was a typo lol.

OpenStudy (michele_laino):

@tom982 I think \cdot is "center dot", namely: \[a*b\]